# Simulation-based inference via telescoping ratio estimation for trawl processes

### v1

@@ Abstract
The growing availability of large and complex datasets has increased interest in temporal stochastic processes that can capture stylized facts such as marginal skewness, non-Gaussian tails, long memory, and even non-Markovian dynamics. While such models are often easy to simulate from, parameter estimation remains challenging. Simulation-based inference (SBI) offers a promising way forward, but existing methods typically require large training datasets or complex architectures and frequently yield confidence (credible) regions that fail to attain their nominal values, raising doubts on the reliability of estimates for the very features that motivate the use of these models. To address these challenges, we propose a fast and accurate, sample-efficient SBI framework for amortized posterior inference applicable to intractable stochastic processes. The proposed approach relies on two main steps: first, we learn the posterior density by decomposing it sequentially across parameter dimensions. Then, we use Chebyshev polynomial approximations to efficiently generate independent posterior samples, enabling accurate inference even when Markov chain Monte Carlo methods mix poorly. We further develop novel diagnostic tools for SBI in this context, as well as post-hoc calibration techniques; the latter not only lead to performance improvements of the learnt inferential tool, but also to the ability to reuse it directly with new time series of varying lengths, thus amortizing the training cost. We demonstrate the method’s effectiveness on trawl processes, a class of flexible infinitely divisible models that generalize univariate Gaussian processes, applied to energy demand data.
@@ 2.1 Lévy bases
and where $\gamma\vcentcolon=\sqrt{\alpha^{2}-\beta^{2}}$ and $K_{\lambda}$ is the modified Bessel function of the second kind of order $\lambda$. This family includes many standard distributions that are supported on the reals as special or limiting cases and offers great flexibility. From a statistical modelling point of view, it is often advantageous to have the marginal law of $L(A)$ in closed form. Many common infinitely divisible distributions (e.g., Poisson, Gaussian, and Gamma) are closed under convolutions, meaning that $L^{\prime}$ and $L(A)$ are from the same named family: for example, if $L^{\prime}\sim\textrm{Poisson}(\nu)$, then $L(A)\sim\textrm{Poisson}(\nu\,\textrm{Leb}(A))$. While the GH family is not convolution-closed in general, it contains two subfamilies that are: the Normal-inverse Gaussian (NIG) distribution, corresponding to $\lambda=-1/2$, and the Variance Gamma (VG) distribution, for $\delta=0$. Indeed, if $L^{\prime}\sim\textrm{NIG}(\alpha,\beta,\delta,\mu)$, then $L(A)\sim\textrm{NIG}\left(\alpha,\beta,\delta\,\mathrm{Leb}\left(A\right),\mu\,\mathrm{Leb}\left(A\right)\right)$ and if $L^{{}^{\prime}}\sim\textrm{VG}(\alpha,\beta,\lambda,\mu)$, then $L(A)\sim\textrm{VG}\left(\alpha,\beta,\lambda\,\mathrm{Leb}(A),\mu\,\mathrm{Leb}(A)\right)$. These classes preserve closed-form marginals while allowing for both exponential and light tails. Parameterizations of all mentioned distributions, along with additional heavy-tailed examples and details on their scaling with $\mathrm{Leb}(A)$, are provided in Section LABEL:suppl_material:parameterizations.
@@ 2.2 Kernel convolutions and trawl processes
The simplest parametric example is the exponential trawl function, where $a(s)=e^{\lambda s}$ for $s\leq 0,\lambda>0$, which yields $\rho(h)=e^{-\lambda h}$ for $h\geq 0$. To interpolate between short and long memory, consider the randomized mixture of exponentials $a(s)=\int_{0}^{\infty}e^{-\lambda s}\pi(\mathrm{d}\lambda)$ for $s\leq 0$, where the decay rate $\lambda$ is sampled according to $\pi$, a probability measure on $\mathbb{R}^{+}$. This recovers the purely exponential case with the Dirac measure $\pi=\delta_{\lambda_{0}}$ for some $\lambda_{0}>0$, and it also includes mixtures of exponentials for $\pi=\sum_{i=1}^{n}\lambda_{i}\delta_{\lambda_{i}}$, with $\sum_{i=1}^{n}\lambda_{i}=1$. In the simulations in Section 4, we use the Inverse Gaussian (IG) trawl function, where $\pi=\textrm{IG}(\gamma,\eta)$ for $\gamma,\,\eta>0$ and $a(s)=\left(1-2s/\gamma^{2}\right)^{-1/2}\exp{\left(\eta\left(1-\sqrt{1-2s/\gamma^{2}}\right)\right)}$, $s\leq 0$, yielding semi-long memory with $\rho(h)=\exp{\left(\eta\left(1-\sqrt{1+2h/\gamma^{2}}\right)\right)}$, $h\geq 0$. By semi-long memory we mean $\rho(\cdot)$ decays slower than exponential but faster than polynomial; see Section LABEL:suppl_material:parameterizations for parameterizations.
@@ 3.2 TRE: a novel divide-and-conquer SBI technique
Figure 1: TRE architectures for learning likelihood ratios. Outputs are on the log scale for stability. (a) Independent classifiers: Each encoder (e.g., Long Short-Term Memory, or LSTM) processes the data (e.g., time series) $\bm{x}$ into summary statistics $s_{i}(\bm{x})$, which are concatenated with $\bm{\theta}$ and passed to a multilayer perceptron (MLP) to approximate $\log{\hat{p}(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})}$. This is added to $\log{p(\theta^{i}\mid\bm{\theta}^{i+1:m})}$ to yield $\log\hat{r}_{i}(\bm{x},\bm{\theta})$. (b) Shared encoder variant: All classifiers share a single encoder with separate MLP heads; see Section LABEL:SM:comparison_with_original_TRE. The $\oplus$ symbol indicates addition by $\log p(\theta^{i}|\bm{\theta}^{i+1:m})$.
@@
See Section LABEL:appendix:subsection_tre for the proof and the general case where $p(\bm{\theta})$ does not factorize. While NRE requires exponentially many samples in $D_{\textrm{KL}}(q_{m}\;\|\;q_{0})$ to accurately represent the BCE loss (Chatterjee & Diaconis 2018), the $i^{\text{th}}$ TRE classifier only requires exponentially many samples in $\mathcal{S}_{i}\vcentcolon=D_{\textrm{KL}}(q_{i}\;\|\;q_{i-1})$. Since all $m$ classifiers can be trained using the same simulated $(\bm{x},\bm{\theta})$ samples, the overall sample complexity scales exponentially in $\max_{1\leq i\leq m}\mathcal{S}_{i}$. Crucially, the order of parameter estimation matters. We assign $\theta^{1}$ to the most influential component and $\theta^{m}$ to the least, since learning $\theta^{m}$ is conditioned on $\bm{\theta}^{1:i-1}$ and is therefore easier, while $\theta^{1}$ must be learned in isolation. This ordering roughly brings the $\mathcal{S}_{i}$ values to similar magnitudes and approximately yields an effective $1/m$ reduction in the exponent. If large discrepancies in $\mathcal{S}_{i}$ appear during training, reordering accordingly improves performance. Note that the terms $\mathcal{S}_{i}=D_{\textrm{KL}}(q_{i}\;\|\;q_{i-1})=\mathbb{E}_{q_{i}(\bm{x},\bm{\theta})}\left[\log{r}_{i}(\bm{x},\bm{\theta})\right]$ can be estimated during training directly from the classifier outputs, with no additional computation. Finally, the $\mathcal{S}_{i}$ values can also inform the allocation of batch sizes across individual classifiers.
@@
We stress that our approach differs fundamentally from Rhodes et al. (2020). While the original TRE therein targets direct density estimation by learning a neural approximation $\hat{p}_{\bm{\psi}}(\cdot)$ of $\bm{x}\mapsto p(\bm{x})$ for a given dataset without an underlying parametric model, our approach is embedded within the SBI framework and learns an amortized approximation $\hat{p}_{\bm{\psi}}(\cdot\mid\cdot)$ of $(\bm{x},\bm{\theta})\mapsto p(\bm{\theta}\mid\bm{x})$. More importantly, the original formulation suffers from a training-inference mismatch due to ratio evaluations in regions of negligible density under the training data, which our approach remedies; see Section LABEL:SM:comparison_with_original_TRE for details.
@@ 3.3 MCMC-free posterior sampling via Chebyshev polynomials
Fast posterior sampling is crucial for both practical applications and the coverage checks from Section 3.5, which require generating posterior samples for thousands of model realizations. Obtaining satisfactory effective sample sizes in this setting, in particular for the simulation study in Section 4, requires adaptive samplers such as NUTS, which tune not only the proposal covariance and step size, but also the number of leapfrog steps. Even this scheme proves inefficient: the adaptive nature prevents any GPU acceleration, makes each chain computationally costly, and forces chains to be run sequentially. As a result, carrying out coverage checks with MCMC is impractical, motivating the development of a novel, MCMC-free approach. However, it is worth noting that when gradient-based MCMC is viable—or when performing maximum likelihood estimation via gradient descent—we can exploit the TRE framework to halve the number of gradient evaluations; see Section LABEL:SM:reduced_gradient_cost.
Having gained access to the one-dimensional conditional densities $\left\{p(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})\right\}_{i=1}^{m}$, we can generate posterior samples sequentially in the components of $\bm{\theta}$. Instead of MCMC, we propose an efficient inversion sampling algorithm, by constructing approximate cumulative distribution functions (CDFs) $\hat{F}\left(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1}\right)$ which are simple to evaluate and then invert by bisection. A detailed account is provided in Section LABEL:suppl:subsection_chebyshev, and we only summarize the main idea here. Following Olver & Townsend (2013), we approximate the one-dimensional densities using Chebyshev polynomials, which only require evaluations at specific Chebyshev knots. This approach has several appealing properties. Computationally, these evaluations only have to be carried out once, regardless of the required number of samples, can be efficiently parallelized on a GPU, do not involve a computationally expensive encoder, and perform reliably in practice. Theoretically, the Chebyshev polynomial approximations converge uniformly on compact intervals and allow both the polynomial and its integral (the approximate CDF) to be evaluated efficiently in closed form, via Clenshaw’s algorithm (Clenshaw 1955), without additional neural network evaluations. Finally, this approach allows us to introduce novel, computationally efficient checks to individually test each component classifier, as explained in Section 3.5. Because the generated samples are independent, we avoid thinning procedures that are often necessary with MCMC samples for simulation-based calibration diagnostics such as rank-based checks (Talts et al. 2018).
@@ 3.4.1 Potential reasons for miscalibration and remedies
where $Y\in\{0,1\}$ denotes the true label of data $\mathbf{Z}$. Intuitively, among all inputs $\mathbf{z}$ for which $c(\mathbf{z})=0.75$, say, about $75\%$ should have label $Y=1$. Nevertheless, neural network-based classifiers, including NRE-based ones, are often miscalibrated (Guo et al. 2017, Hermans et al. 2022). Miscalibration is typically quantified using the expected calibration error (ECE), $\mathbb{E}\left[\left|\mathds{P}\left(Y=1\mid c(\mathbf{Z})\right)-c(\mathbf{Z})\right|\right]$, which measures the average discrepancy between the predicted confidence $c(\mathbf{Z})$ and the accuracy $\mathds{P}\left(Y=1\mid c(\mathbf{Z})\right)$, and can be approximated from samples as discussed in Section LABEL:SM:ECE. Empirical studies (Niculescu-Mizil & Caruana 2005, Guo et al. 2017, Minderer et al. 2021) report conflicting results, with calibration varying based on regularization, architecture and model capacity. We identify two main avenues for improvement and further leverage the rich literature on post-hoc calibration of probabilistic classifiers to ensure that the level curves of the likelihood (posterior) are calibrated.
@@ 4.1 Simulation setting
We apply our methodology to a class of trawl processes $X$ capable of exhibiting short and long memory, semi-heavy tails and skewness. To this end, consider $X_{t}\sim\textrm{NIG}(\mu,\sigma,\beta)$ with autocorrelation function $\rho(h)=\exp{\left(\,\eta_{\textrm{acf}}\left(1-\sqrt{1+2h/\gamma_{\textrm{acf}}^{2}}\right)\right)}$ for $h\geq 0$. We adopt a three-parameter form for the NIG distribution, see Section LABEL:subsubsection:alternative_specification, instead of the conventional four-parameter form for two reasons. First, under the standard four-parameter formulation, distinct parameter values can yield densities that are visually indistinguishable, hindering inference and posterior sampling. Second, our reparameterization remains interpretable, as $\mu$ and $\sigma$ correspond to the mean and standard deviation. We set independent uniform sampling distributions to ensure the amortized model is valid for a wide range of cases: $\gamma_{\text{acf}}\sim\mathcal{U}(10,20),\,\eta_{\text{acf}}\sim\mathcal{U}(10,20),\,\mu\sim\mathcal{U}(-1,1),\,\sigma\sim\mathcal{U}(0.5,1.5),\,\beta\sim\mathcal{U}(-5,5)$, and let $\bm{\theta}=\left(\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu,\sigma,\beta\right)$. The support restrictions on $\mu$ and $\sigma$ are without loss of generality, as we can feed the classifier time series that are centered and scaled with the empirical mean and standard deviation, and invert the transformation afterwards.
As discussed in Section 3.2, coordinate ordering matters in TRE. The autocorrelation functions (ACF) parameters are weakly identifiable, hence we infer these jointly. We thus learn a block of two coordinates, and we place it first in the TRE decomposition below to reduce the computational cost of building the Chebyshev approximation; see Section LABEL:SM:seq_p_sam. We then order the remaining parameters by expected learning difficulty: mean, standard deviation, then tilt. Formally, let $\bm{x}$ be an observation of $X$ at times $1,\ldots,k=1500$. Then, the ratio $r(\bm{x},\bm{\theta})$ can be learnt by training four classifiers:
@@
We monitor the BCE loss, cross-entropy $\mathcal{S}$, accuracy, and balancing metric $\mathcal{B}$ (see Section 3.5.3) throughout training and display the results in Figure 2. All metrics approximately stabilize, suggesting convergence. Gradients are computed with data simulated on-the-fly at each training iteration, whereas the metrics are evaluated on a fixed, holdout dataset. We use the architecture from Figure 1(a), in which trawl process realizations $\bm{x}$ are fed through an LSTM encoder. The encoded features are concatenated with the relevant parameter subset of $\bm{\theta}$ and then passed through fully connected layers; implementation details are available in Section LABEL:section:training_details.

### v2

@@ Abstract
The growing availability of large and complex datasets has increased interest in temporal stochastic processes that can capture stylized facts such as marginal skewness, non-Gaussian tails, long memory, and even non-Markovian dynamics. While such models are often easy to simulate from, parameter estimation remains challenging. Simulation-based inference (SBI) offers a promising way forward, but existing methods typically require large training datasets or complex architectures and frequently yield confidence (credible) regions that fail to attain their nominal values, raising doubts on the reliability of estimates for the very features that motivate the use of these models. To address these challenges, we propose a fast and accurate, sample-efficient SBI framework for amortized posterior inference applicable to intractable stochastic processes. The proposed approach relies on two main steps: first, we learn the posterior density by decomposing it sequentially across parameter dimensions. Then, we use Chebyshev polynomial approximations to efficiently generate independent posterior samples, enabling accurate inference even when Markov chain Monte Carlo methods mix poorly. We further develop novel diagnostic tools for SBI in this context, as well as post-hoc calibration techniques; the latter not only lead to performance improvements of the learned inferential tool, but also to the ability to reuse it directly with new time series of varying lengths, thus amortizing the training cost. We demonstrate the method’s effectiveness on trawl processes, a class of flexible infinitely divisible models that generalize univariate Gaussian processes, applied to energy demand data.
@@ 2.1 Lévy bases
and where $\gamma\vcentcolon=\sqrt{\alpha^{2}-\beta^{2}}$ and $K_{\lambda}$ is the modified Bessel function of the second kind of order $\lambda$. This family includes many standard distributions that are supported on the reals as special or limiting cases and offers great flexibility. From a statistical modelling point of view, it is often advantageous to have the marginal law of $L(A)$ in closed form. Many common infinitely divisible distributions (e.g., Poisson, Gaussian, and Gamma) are closed under convolutions, meaning that $L^{\prime}$ and $L(A)$ are from the same named family: for example, if $L^{\prime}\sim\textrm{Poisson}(\nu)$, then $L(A)\sim\textrm{Poisson}(\nu\,\textrm{Leb}(A))$. While the GH family is not convolution-closed in general, it contains two subfamilies that are: the Normal-inverse Gaussian (NIG) distribution, corresponding to $\lambda=-1/2$, and the Variance Gamma (VG) distribution, for $\delta=0$. Indeed, if $L^{\prime}\sim\textrm{NIG}(\alpha,\beta,\delta,\mu)$, then $L(A)\sim\textrm{NIG}\left(\alpha,\beta,\delta\,\mathrm{Leb}\left(A\right),\mu\,\mathrm{Leb}\left(A\right)\right)$ and if $L^{{}^{\prime}}\sim\textrm{VG}(\alpha,\beta,\lambda,\mu)$, then $L(A)\sim\textrm{VG}\left(\alpha,\beta,\lambda\,\mathrm{Leb}(A),\mu\,\mathrm{Leb}(A)\right)$. These classes preserve closed-form marginals while allowing for both exponential and light tails. Parameterizations of all mentioned distributions, along with additional heavy-tailed examples and details on their scaling with $\mathrm{Leb}(A)$, are provided in Section S1.
@@ 2.2 Kernel convolutions and trawl processes
The simplest parametric example is the exponential trawl function, where $a(s)=e^{\lambda s}$ for $s\leq 0,\lambda>0$, which yields $\rho(h)=e^{-\lambda h}$ for $h\geq 0$. To interpolate between short and long memory, consider the randomized mixture of exponentials $a(s)=\int_{0}^{\infty}e^{-\lambda s}\pi(\mathrm{d}\lambda)$ for $s\leq 0$, where the decay rate $\lambda$ is sampled according to $\pi$, a probability measure on $\mathbb{R}^{+}$. This recovers the purely exponential case with the Dirac measure $\pi=\delta_{\lambda_{0}}$ for some $\lambda_{0}>0$, and it also includes mixtures of exponentials for $\pi=\sum_{i=1}^{n}\lambda_{i}\delta_{\lambda_{i}}$, with $\sum_{i=1}^{n}\lambda_{i}=1$. In the simulations in Section 4, we use the Inverse Gaussian (IG) trawl function, where $\pi=\textrm{IG}(\gamma,\eta)$ for $\gamma,\,\eta>0$ and $a(s)=\left(1-2s/\gamma^{2}\right)^{-1/2}\exp{\left(\eta\left(1-\sqrt{1-2s/\gamma^{2}}\right)\right)}$, $s\leq 0$, yielding semi-long memory with $\rho(h)=\exp{\left(\eta\left(1-\sqrt{1+2h/\gamma^{2}}\right)\right)}$, $h\geq 0$. By semi-long memory we mean $\rho(\cdot)$ decays slower than exponential but faster than polynomial; see Section S1 for parameterizations.
@@ 3.2 TRE: a novel divide-and-conquer SBI technique
Figure 1: TRE architectures for learning likelihood ratios. Outputs are on the log scale for stability. (a) Independent classifiers: Each encoder (e.g., Long Short-Term Memory, or LSTM) processes the data (e.g., time series) $\bm{x}$ into summary statistics $s_{i}(\bm{x})$, which are concatenated with $\bm{\theta}$ and passed to a multilayer perceptron (MLP) to approximate $\log{\hat{p}(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})}$. This is added to $\log{p(\theta^{i}\mid\bm{\theta}^{i+1:m})}$ to yield $\log\hat{r}_{i}(\bm{x},\bm{\theta})$. (b) Shared encoder variant: All classifiers share a single encoder with separate MLP heads; see Section S2.3. The $\oplus$ symbol indicates addition by $\log p(\theta^{i}|\bm{\theta}^{i+1:m})$.
@@
See Section S2 for the proof and the general case where $p(\bm{\theta})$ does not factorize. While NRE requires exponentially many samples in $D_{\textrm{KL}}(q_{m}\;\|\;q_{0})$ to accurately represent the BCE loss (Chatterjee & Diaconis 2018), the $i^{\text{th}}$ TRE classifier only requires exponentially many samples in $\mathcal{S}_{i}\vcentcolon=D_{\textrm{KL}}(q_{i}\;\|\;q_{i-1})$. Since all $m$ classifiers can be trained using the same simulated $(\bm{x},\bm{\theta})$ samples, the overall sample complexity scales exponentially in $\max_{1\leq i\leq m}\mathcal{S}_{i}$. Crucially, the order of parameter estimation matters. We assign $\theta^{1}$ to the most influential component and $\theta^{m}$ to the least, since learning $\theta^{m}$ is conditioned on $\bm{\theta}^{1:i-1}$ and is therefore easier, while $\theta^{1}$ must be learned in isolation. This ordering roughly brings the $\mathcal{S}_{i}$ values to similar magnitudes and approximately yields an effective $1/m$ reduction in the exponent. If large discrepancies in $\mathcal{S}_{i}$ appear during training, reordering accordingly improves performance. Note that the terms $\mathcal{S}_{i}=D_{\textrm{KL}}(q_{i}\;\|\;q_{i-1})=\mathbb{E}_{q_{i}(\bm{x},\bm{\theta})}\left[\log{r}_{i}(\bm{x},\bm{\theta})\right]$ can be estimated during training directly from the classifier outputs, with no additional computation. Finally, the $\mathcal{S}_{i}$ values can also inform the allocation of batch sizes across individual classifiers.
@@
We stress that our approach differs fundamentally from Rhodes et al. (2020). While the original TRE therein targets direct density estimation by learning a neural approximation $\hat{p}_{\bm{\psi}}(\cdot)$ of $\bm{x}\mapsto p(\bm{x})$ for a given dataset without an underlying parametric model, our approach is embedded within the SBI framework and learns an amortized approximation $\hat{p}_{\bm{\psi}}(\cdot\mid\cdot)$ of $(\bm{x},\bm{\theta})\mapsto p(\bm{\theta}\mid\bm{x})$. More importantly, the original formulation suffers from a training-inference mismatch due to ratio evaluations in regions of negligible density under the training data, which our approach remedies; see Section S2.3 for details.
@@ 3.3 MCMC-free posterior sampling via Chebyshev polynomials
Fast posterior sampling is crucial for both practical applications and the coverage checks from Section 3.5, which require generating posterior samples for thousands of model realizations. Obtaining satisfactory effective sample sizes in this setting, in particular for the simulation study in Section 4, requires adaptive samplers such as NUTS, which tune not only the proposal covariance and step size, but also the number of leapfrog steps. Even this scheme proves inefficient: the adaptive nature prevents any GPU acceleration, makes each chain computationally costly, and forces chains to be run sequentially. As a result, carrying out coverage checks with MCMC is impractical, motivating the development of a novel, MCMC-free approach. However, it is worth noting that when gradient-based MCMC is viable—or when performing maximum likelihood estimation via gradient descent—we can exploit the TRE framework to halve the number of gradient evaluations; see Section S2.4.
Having gained access to the one-dimensional conditional densities $\left\{p(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})\right\}_{i=1}^{m}$, we can generate posterior samples sequentially in the components of $\bm{\theta}$. Instead of MCMC, we propose an efficient inversion sampling algorithm, by constructing approximate cumulative distribution functions (CDFs) $\hat{F}\left(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1}\right)$ which are simple to evaluate and then invert by bisection. A detailed account is provided in Section S3, and we only summarize the main idea here. Following Olver & Townsend (2013), we approximate the one-dimensional densities using Chebyshev polynomials, which only require evaluations at specific Chebyshev knots. This approach has several appealing properties. Computationally, these evaluations only have to be carried out once, regardless of the required number of samples, can be efficiently parallelized on a GPU, do not involve a computationally expensive encoder, and perform reliably in practice. Theoretically, the Chebyshev polynomial approximations converge uniformly on compact intervals and allow both the polynomial and its integral (the approximate CDF) to be evaluated efficiently in closed form, via Clenshaw’s algorithm (Clenshaw 1955), without additional neural network evaluations. Finally, this approach allows us to introduce novel, computationally efficient checks to individually test each component classifier, as explained in Section 3.5. Because the generated samples are independent, we avoid thinning procedures that are often necessary with MCMC samples for simulation-based calibration diagnostics such as rank-based checks (Talts et al. 2018).
@@ 3.4.1 Potential reasons for miscalibration and remedies
where $Y\in\{0,1\}$ denotes the true label of data $\mathbf{Z}$. Intuitively, among all inputs $\mathbf{z}$ for which $c(\mathbf{z})=0.75$, say, about $75\%$ should have label $Y=1$. Nevertheless, neural network-based classifiers, including NRE-based ones, are often miscalibrated (Guo et al. 2017, Hermans et al. 2022). Miscalibration is typically quantified using the expected calibration error (ECE), $\mathbb{E}\left[\left|\mathds{P}\left(Y=1\mid c(\mathbf{Z})\right)-c(\mathbf{Z})\right|\right]$, which measures the average discrepancy between the predicted confidence $c(\mathbf{Z})$ and the accuracy $\mathds{P}\left(Y=1\mid c(\mathbf{Z})\right)$, and can be approximated from samples as discussed in Section S4.1. Empirical studies (Niculescu-Mizil & Caruana 2005, Guo et al. 2017, Minderer et al. 2021) report conflicting results, with calibration varying based on regularization, architecture and model capacity. We identify two main avenues for improvement and further leverage the rich literature on post-hoc calibration of probabilistic classifiers to ensure that the level curves of the likelihood (posterior) are calibrated.
@@ 4.1 Simulation setting
We apply our methodology to a class of trawl processes $X$ capable of exhibiting short and long memory, semi-heavy tails and skewness. To this end, consider $X_{t}\sim\textrm{NIG}(\mu,\sigma,\beta)$ with autocorrelation function $\rho(h)=\exp{\left(\,\eta_{\textrm{acf}}\left(1-\sqrt{1+2h/\gamma_{\textrm{acf}}^{2}}\right)\right)}$ for $h\geq 0$. We adopt a three-parameter form for the NIG distribution, see Section S1.2, instead of the conventional four-parameter form for two reasons. First, under the standard four-parameter formulation, distinct parameter values can yield densities that are visually indistinguishable, hindering inference and posterior sampling. Second, our reparameterization remains interpretable, as $\mu$ and $\sigma$ correspond to the mean and standard deviation. We set independent uniform sampling distributions to ensure the amortized model is valid for a wide range of cases: $\gamma_{\text{acf}}\sim\mathcal{U}(10,20),\,\eta_{\text{acf}}\sim\mathcal{U}(10,20),\,\mu\sim\mathcal{U}(-1,1),\,\sigma\sim\mathcal{U}(0.5,1.5),\,\beta\sim\mathcal{U}(-5,5)$, and let $\bm{\theta}=\left(\gamma_{\textrm{acf}},\eta_{\textrm{acf}},\mu,\sigma,\beta\right)$. The support restrictions on $\mu$ and $\sigma$ are without loss of generality, as we can feed the classifier time series that are centered and scaled with the empirical mean and standard deviation, and invert the transformation afterwards.
As discussed in Section 3.2, coordinate ordering matters in TRE. The autocorrelation functions (ACF) parameters are weakly identifiable, hence we infer these jointly. We thus learn a block of two coordinates, and we place it first in the TRE decomposition below to reduce the computational cost of building the Chebyshev approximation; see Section S3.4. We then order the remaining parameters by expected learning difficulty: mean, standard deviation, then tilt. Formally, let $\bm{x}$ be an observation of $X$ at times $1,\ldots,k=1500$. Then, the ratio $r(\bm{x},\bm{\theta})$ can be learnt by training four classifiers:
@@
We monitor the BCE loss, cross-entropy $\mathcal{S}$, accuracy, and balancing metric $\mathcal{B}$ (see Section 3.5.3) throughout training and display the results in Figure 2. All metrics approximately stabilize, suggesting convergence. Gradients are computed with data simulated on-the-fly at each training iteration, whereas the metrics are evaluated on a fixed, holdout dataset. We use the architecture from Figure 1(a), in which trawl process realizations $\bm{x}$ are fed through an LSTM encoder. The encoded features are concatenated with the relevant parameter subset of $\bm{\theta}$ and then passed through fully connected layers; implementation details are available in Section S5.3.